OpenStudy (anonymous):

For which value of x does the graph of y = 2x2 − 7x + 6 cross the x-axis? is it 2

OpenStudy (anonymous):

at x axis,y=0 so, put y=0 0=2x2-7x+6 solving quadratic 2x2-4x-3x+6 2x(x-2)-3(x-2)=0 (2x-3)(x-2)=0 x=3/2,2

OpenStudy (anonymous):

So it is 2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can you help me with one more

OpenStudy (anonymous):

Of course

OpenStudy (anonymous):

What are the approximate solutions of 4x2 + 3 = −12x to the nearest hundredth? x ≈ −3.23 and x ≈ 0.23 x ≈ −2.72 and x ≈ −0.28 x ≈ 0.28 and x ≈ 2.72 x ≈ −0.23 and x ≈ 3.23

OpenStudy (anonymous):

I got the second choice

OpenStudy (anonymous):

To solve a quadratic equation, put the trinomial on one side and make it equal to 0. Then factor it and make each set of parentheses equal to 0. Then solve both equations to get your solution. 4x² + 3 = -12x +12x = +12x 4x² + 12x + 3 = 0 x = -0.28 , -2.73

OpenStudy (anonymous):

So Which answer would it be @ziko1995

OpenStudy (anonymous):

The second one :D

OpenStudy (anonymous):

Ok haha